Three identical metallic conducting spheres carry the following charges: q11 = +7.00 μC, q22 = +7.00 μC, and q33 = −2.80 μC. The spheres that carry the charges q11 and q22 are brought into contact. Then they are separated. After that, one of those two spheres is brought into contact with the third sphere that carries the charge q33; those two are then separated as well.
a) What is the final charge on the third sphere?

b) How many excess (or deficiency) electrons make up the final charge on the third sphere? (answer in microcoulombs)
Answer #1

Given q_{1}=7mu C , q_{2}=7mu C and q_{3}=-2.8mu C

Since charge on q1 and q2 are same ,when they are brought in contact their charges remains same

Charge on q3 is different than other 2

If either one of first two spheres are brought in contact with q3 ,the charge should equalize

(q_{1}+q_{3})/2 =(7mu C-2.8mu C)/2 =2.1mu C

(a)ANSWER: Final charge on {color{Red} q_{3}=2.1mu C}

===================

Since charge is positive electrons are deficient

No of electrons=?

No of electrons*charge of 1 electron = charge on q3

n*e=q_{3}

n * 1.602 * 10-19C-2.1 * 10-0C

n = 2.1 * 10-0/1.602 * 10-1

ANSWER: n 1.31 * 1013

========================